# Rank, Chunk, and Expand: Lineage-Oriented Reasoning for Taxonomy Expansion

### v1

@@ Abstract
Taxonomies are hierarchical knowledge graphs crucial for recommendation systems, and web applications. As data grows, expanding taxonomies is essential, but existing methods face key challenges: (1) discriminative models struggle with representation limits and generalization, while (2) generative methods either process all candidates at once, introducing noise and exceeding context limits, or discard relevant entities by selecting noisy candidates. We propose LORex (Lineage-Oriented Reasoning for Taxonomy Expansion), a plug-and-play framework that combines discriminative ranking and generative reasoning for efficient taxonomy expansion. Unlike prior methods, LORex ranks and chunks candidate terms into batches, filtering noise and iteratively refining selections by reasoning candidates’ hierarchy to ensure contextual efficiency. Extensive experiments across four benchmarks and twelve baselines show that LORex improves accuracy by 12% and Wu & Palmer similarity by 5% over state-of-the-art methods.
Rank, Chunk, and Expand: Lineage-Oriented Reasoning for Taxonomy Expansion
@@ 1 Introduction
Taxonomies are hierarchical graph structures that capture hypernymy (“is-a”) relationships among concepts, making them essential for knowledge organization across various domains. Conglomerates leverage them to power search engines Strzelecki and Rutecka (2019), enhance product recommendations Mao et al. (2020); Karamanolakis et al. (2020) and improve advertisements Gonçalves et al. (2019); Manzoor et al. (2020). Despite their utility, real-world taxonomies are manually created by domain experts, which is cumbersome and costly, limiting their ability to capture emerging concepts Jurgens and Pilehvar (2016). Additionally, as new concepts continuously emerge, updating taxonomies manually becomes impractical.
@@
To overcome the limitations of manual taxonomy construction, research has now shifted toward taxonomy expansion, which involves the integration of new entities into an existing seed taxonomy Jurgens and Pilehvar (2016); Bordea et al. (2016a). This process positions new concepts under appropriate existing nodes, known as anchor nodes, to preserve the structural integrity of the original taxonomy. As illustrated in Figure 1, this ensures that new terms, like “Marine Pollution”, are inserted under “Water Pollution” rather than the broader “Pollution”, maintaining hierarchical coherence.
@@
The current generation research on taxonomy expansion leverages LLMs to integrate new concepts into existing taxonomies. Models like LLaMA Dubey et al. (2024) and GPT-4 Achiam et al. (2023), excel in structural knowledge comprehension Wei et al. (2023); Liu et al. (2025). Notably, GPT-4, with 1.7 trillion parameters, generates parent terms directly Zeng et al. (2024b, a), while LLaMA-2 (7B parameters) requires fine-tuning on self-supervised annotation data for taxonomy expansion Mishra et al. (2024); Xu et al. (2025).
@@
We conduct experiments to evaluate TEMPORA for candidate ranking and LORex for taxonomy expansion on four public benchmark datasets, including SemEval-2016 Task 13 Taxonomy Extraction Evaluation Bordea et al. (2016a) and WordNet Bansal et al. (2014). TEMPORA outperforms all ranking baselines, improving Hit@k by 21.3%, despite their extensive self-supervised training. LORex surpasses eight baselines across four benchmarks, achieving 12% higher accuracy and a 5% gain in Wu & Palmer (Wu&P) metric. Furthermore, we perform ablations to analyze the performance of LORex under varying conditions, including its effectiveness with different open-source LLMs.
We summarize our main contributions below: [^1] Source code: https://github.com/sahilmishra0012/lorex.
@@
Figure 2: An illustration of our proposed framework – LORex, which first ranks candidates using TEMPORA and chunks them. The semantic filter module then checks the batch relevance, which further leads to parent retrieval and verification. In Euler tour, $\bullet$ ($\blacklozenge$) means “is child of” (“is parent of”) relationship.
@@ 2 Related Work
Taxonomy Expansion. Expert-curated taxonomies, such as The Plant List Schellenberger Costa et al. (2023) and NCBI Taxonomy Schoch et al. (2020), require continuous updates, necessitating robust taxonomy expansion. Recent discriminative methods leverage structural features such as taxonomy paths Liu et al. (2021), mini-paths Yu et al. (2020), hierarchical neighbourhoodJiang et al. (2022), ego-trees Wang et al. (2021), and ego networks Shen et al. (2020) to enhance anchor node representation and model hypernymy relationships. Prompting frameworks such as TaxoPrompt Xu et al. (2022), ATTEMPT Xia et al. (2023), and TacoPrompt Xu et al. (2023) model taxonomy paths via masked language modeling. However, these methods struggle with limited self-supervised annotated data and small-scale encoders. To address this, LLM-based solutions such as FLAME Mishra et al. (2024) and COMI Xu et al. (2025) fine-tune LLaMA models, while CodeTaxo Zeng et al. (2024b) employs GPT-4 inference. Though effective, fine-tuning is computationally intensive, and GPT-4 API usage is costly. To address these challenges, our proposed LORex model integrates discriminative ranking with generative reasoning, leveraging smaller language models. LORex employs iterative prompting to systematically reason over taxonomic lineages, ensuring computational efficiency while maintaining high accuracy in taxonomy expansion.
@@ 4 Proposed Method
LORex, a plug-and-play taxonomy expansion framework (Figure 2), integrates discriminative ranking (TEMPORA) with iterative structure reasoning to integrate new concepts into an existing taxonomy.
@@ 4.1 TEMPORA
To effectively capture the hierarchical structure, we extract the path from the root node to the anchor node, along with its children and siblings. This hierarchy is then verbalized by performing an Euler tour on the tree, which begins at the anchor node, ascends to the root, and subsequently traverses siblings, followed by child nodes in a structured manner. The Euler-toured path is represented as $P=[n_{p},n_{p-1},\allowbreak\dots,n_{o},n_{1},\dots,\allowbreak n_{p-1},n_{s_{1}},\dots,n_{p-1},n_{p},n_{c_{1}},n_{p},n_{c_{2}}]$, where $n_{o}$ is the root node while $n_{p-i}$, $n_{s_{i}}$ and $n_{c_{i}}$ are the ancestor siblings and children of anchor node, respectively. Unlike TEMP, we use interpretable relational phrases such as “is parent of” and “is child of” to enhance the clarity and interpretability of the Euler-toured path, returning verbalized path as $P_{v}$. We return two different verbalized paths $P_{v}(\!q,\!P)\!=\!\textit{Verbalizer}(\!q,\!P)$ and $P_{\!v\!}(P)\!=\!\textit{Verbalizer}(\!P)$ where $q$ is the query node whose definition is used by the verbalizer, which is concatenated with path $P$ via the ([SEP]) token. The pre-trained encoder returns the following vectors,
@@ 4.4 Parent Verifier Module
In order to verify if the retrieved parent from LLM is the most appropriate one, existing works have employed discriminative verifiers Zeng et al. (2024a). However, these verifiers are often limited by their representation power. Moreover, if we use path-based verifier similar to TEMPORA, the verifier is again biased towards the ranked candidates, creating a bottleneck in verification. To address this, we leverage LLMs to rank candidate paths in the batch. Specifically, we utilize token probabilities to rank candidate paths.
The LLM is prompted with the hierarchy of retrieved candidates along with the remaining candidates within the batch to select the most appropriate path (see Appendix D for the prompt). Then, average token log probabilities are computed for all the paths in the batch to rank them to eliminate the possibility of the verifier selecting longer paths. The highest-scored candidate path is computed as,
@@ 5.1 Experimental Setup
We experiment on four open-source benchmarks: three from SemEval-2016 Task 13 Taxonomy Extraction Evaluation Bordea et al. (2016b) – Science (SemEval-Sci or Sci), Environment (SemEval-Env or Env) and Food (SemEval-Food) while the other is a collection of 114 WordNet sub-taxonomies (with 10 to 50 nodes each) Bansal et al. (2014) (refer to Table 6 for statistics on benchmarks and Appendix G for train-test split and procedure to resolve definitions).
@@ 5.2 Performance Comparison
Table 2 shows the performance of LORex on four benchmarks. As discussed in Appendices C and D, our framework uses prompts in an instruction-tuned paradigm. Therefore, we three instruction-tuned LLMs – LLaMA-3.1-8B-Instruct (LORex8B-3.1I), LLaMA-3-8B-Instruct (LORex8B-3I), and LLaMA-2-7B-Chat (LORex7B-2C) to compare against state-of-the-art methods. As shown, LORex surpasses all first and second-generation discriminative and third-generation prompting methods, with LORex8B-3.1I achieving the best performance. Specifically, LORex achieves 12% improvement in accuracy and 5% improvement in Wu&P over the best extensively trained generative baseline, FLAME.
@@ 5.3 Ablation Studies
We conduct ablations to analyze every component of LORex. Specifically, we analyze (i) the impact of hierarchical and semantic reasoning on parent retrieval, (ii) effect of different types of reasoning verifiers, (iii) the influence of discriminative ranker, and (iv) varying chunk sizes on the overall performance of LORex.
@@
We analyze the impact of hierarchical and semantic reasoning on parent retrieval in LORex, as shown in Figure 3. Using both paths and definitions gives the highest performance while removing both components leads to the worst performance. Removing definitions reduces accuracy by 17.50% and Wu&P by 2.78%. Further, removing hierarchical paths exacerbates the performance degradation, leading to a 31.99% decrease in accuracy and a 9.78% reduction in Wu&P scores.
Figure 3: Performance comparison of hierarchical and semantic reasoning on parent retrieval across all benchmarks (W/O: without).
@@
Figure 4 compares LORex’s performance under different verification strategies. The path-based verifier achieves the highest accuracy, outperforming the candidate-based verifier (CBV) by 47.04% in accuracy and 10.84% in Wu&P scores. Even without a verifier (WoV), retrieval remains 21.31% more accurate and 7.76% higher in Wu&P than CBV. The lowest performance is observed with CBV, where path information is absent, and verification relies solely on candidate evaluation, which introduces noise during verification and rejects relevant candidates. These results emphasize the importance of hierarchical reasoning during verification.
Query Parent Retriever Path-Based Verifier Candidate-Based Verifier
@@
Figure 4: Performance comparison of verifier variants on SemEval-2016. "WoV" = without verifier, "CBV" = candidate-based verifier, "PBV" = path-based verifier.
Method Env Sci WordNet
Acc Wu&P Acc Wu&P Acc Wu&P
TEMPORA 67.31 82.99 64.71 87.42 49.78 84.58
Random Shuffle 5.76 47.64 2.35 48.95 1.51 38.95
Impact of Discriminative Ranker LORex’s Performance.
Table 4 compares LORex’s performance with a discriminative ranker versus random candidate shuffling. The discriminative ranker improves accuracy by 57.33% and Wu&P scores by 39.8%, highlighting its role in filtering relevant candidates and enhancing retrieval. Without structured ranking, the retriever and verifier struggle to select relevant candidates, as random shuffling introduces noise by placing semantically similar but incorrect candidates before the relevant ones, which prevents moving to chunks containing relevant candidates.
@@
We analyze the impact of varying chunk sizes on LORex’s retrieval performance across SemEval-2016 benchmarks, as shown in Figure 5. The results indicate that smaller chunk sizes (2–3) lead to performance degradation due to limited contextual information, while excessively large chunk sizes ($\geq$20) reduce effectiveness as the retriever encounters an increased number of irrelevant candidates, introducing noise in the chunk. The best performance is achieved with chunk sizes between 4 and 10, indicating a balance between relevant context depth and noise.

### v4

@@ Abstract
Taxonomies are hierarchical knowledge graphs crucial for search engines, recommendation systems, and web applications. As data grows, expanding taxonomies is essential, but existing methods face key challenges: (1) discriminative models struggle with representation limits and generalization, while (2) generative methods either process all candidates at once, introducing noise and exceeding context limits, or discard relevant entities by selecting a sample of candidates. We propose LORex (Lineage-Oriented Reasoning for Taxonomy Expansion), a plug-and-play framework that combines discriminative ranking and generative reasoning for efficient taxonomy expansion without fine-tuning. Unlike prior methods, LORex ranks and chunks candidate terms into batches, filtering noise and iteratively refining selections by reasoning candidates’ hierarchy to ensure contextual efficiency. Extensive experiments across four benchmarks and twelve baselines show that LORex improves accuracy by 12% and Wu & Palmer similarity 5% over state-of-the-art methods.
Rank, Chunk, and Expand:
Lineage-Oriented Reasoning for Taxonomy Expansion
@@ 1 Introduction
Taxonomies are hierarchical graph structures that capture hypernymy (“is-a”) relationships among concepts, making them essential for knowledge organization across various domains. Conglomerates leverage them to power search engines Strzelecki and Rutecka (2019); Janssen and Proper (2021), enhance product recommendations Mao et al. (2020); Karamanolakis et al. (2020) and improve advertisements Gonçalves et al. (2019); Manzoor et al. (2020). Despite their utility, real-world taxonomies are manually created by domain experts, which is cumbersome and costly, limiting their ability to capture emerging concepts Jurgens and Pilehvar (2016); Bordea et al. (2016a). Additionally, as new concepts continuously emerge, updating taxonomies manually becomes increasingly impractical.
@@
To overcome the limitations of manual taxonomy construction, research has now shifted toward taxonomy expansion, which involves the integration of new entities into an existing seed taxonomy Jurgens and Pilehvar (2016); Bordea et al. (2016a). This process positions new concepts under appropriate existing nodes, known as anchor nodes, to preserve the structural integrity of the original taxonomy. As illustrated in Fig. 1, this ensures that new terms, like “Marine Pollution”, are inserted under “Water Pollution” rather than the broader “Pollution”, maintaining hierarchical coherence.
@@
The current generation research on taxonomy expansion leverages LLMs to integrate new concepts into existing taxonomies. Models like LLaMA Touvron et al. (2023); Dubey et al. (2024) and GPT-4 Achiam et al. (2023), excel in structural knowledge comprehension Wei et al. (2023); Liu et al. (2025). Notably, GPT-4, with 1.7 trillion parameters, generates parent terms directly Zeng et al. (2024b, a), while LLaMA-2 (7B parameters) requires fine-tuning on self-supervised annotation data for taxonomy expansion Mishra et al. (2024); Xu et al. (2025).
@@
We conduct extensive experiments to evaluate TEMPORA for candidate ranking and LORex for taxonomy expansion on four public benchmark datasets, including SemEval-2016 Task 13 Taxonomy Extraction Evaluation Bordea et al. (2016a) and WordNet Bansal et al. (2014). TEMPORA outperforms all ranking baselines, improving Hit@k by 21.3%, despite their extensive self-supervised training. LORex surpasses eight baselines across four benchmarks, achieving 12% higher accuracy and a 5% gain in Wu & Palmer (Wu&P) metric. Furthermore, we perform ablations to analyze the performance of LORex under varying conditions, including its effectiveness with different open-source LLMs.
We summarize our main contributions below: [^1] Code: https://github.com/sahilmishra0012/LOREx.
@@
Figure 2: An illustration of our proposed framework, LORex, which first ranks candidates using TEMPORA and chunks them. The semantic filter module then checks the batch relevance, which further leads to parent retrieval and verification. In Euler tour, $\bullet$ ($\blacklozenge$) means “is child of” (“is parent of”) relationship.
@@ 2 Related Work
Taxonomy Expansion. Expert-curated taxonomies, such as The Plant List Schellenberger Costa et al. (2023) and NCBI Taxonomy Schoch et al. (2020), require continuous updates, necessitating robust taxonomy expansion. Recent discriminative methods leverage structural features such as taxonomy paths Liu et al. (2021), mini-paths Yu et al. (2020), hierarchical neighbourhoodJiang et al. (2022), ego-trees Wang et al. (2021), ego networks Shen et al. (2020) and even quantum entanglements Mishra et al. (2025), to enhance anchor node representation and model hypernymy relationships. Prompting frameworks such as TaxoPrompt Xu et al. (2022), ATTEMPT Xia et al. (2023), and TacoPrompt Xu et al. (2023) model taxonomy paths via masked language modeling. However, these methods struggle with limited self-supervised annotated data and small-scale encoders. To address this, LLM-based solutions such as FLAME Mishra et al. (2024) and COMI Xu et al. (2025) fine-tune LLaMA models, while CodeTaxo Zeng et al. (2024b) employs GPT-4 inference. Though effective, fine-tuning is computationally intensive, and GPT-4 API usage is costly. To address these challenges, our proposed LORex model integrates discriminative ranking with generative reasoning, leveraging smaller language models. LORex employs iterative prompting to systematically reason over taxonomic lineages, ensuring computational efficiency while maintaining high accuracy in taxonomy expansion.
@@ 3 Problem Definition
As new entities continue to emerge, the key challenge is incorporating them into an existing taxonomy $\mathcal{T}^{o}$, referred to as taxonomy expansion,
@@ 4 Proposed Method
LORex, a plug-and-play taxonomy expansion framework (Fig. 2), integrates discriminative ranking (TEMPORA) with iterative structure reasoning to integrate new concepts into an existing taxonomy.
@@ 4.1 TEMPORA
To effectively capture the hierarchical structure, we extract the path from the root node to the anchor node, along with its children and siblings. This hierarchy is then verbalized by performing an Euler tour on the tree, which begins at the anchor node, ascends to the root, and subsequently traverses siblings, followed by child nodes in a structured manner. The Euler-toured path is represented as $P=[n_{p},n_{p-1},\allowbreak\dots,n_{o},n_{1},\dots,\allowbreak n_{p-1},n_{s_{1}},\dots,n_{p-1},n_{p},n_{c_{1}},n_{p},n_{c_{2}}]$, where $n_{o}$ is the root node while $n_{p-i}$, $n_{s_{i}}$ and $n_{c_{i}}$ are the ancestor siblings and children of anchor node, respectively. Unlike traditional Euler tour, where each node can only be visited once, our version of Euler tour can visit a node more than once. Contrary to TEMP, we use interpretable relational phrases such as “is parent of” and “is child of” to enhance the clarity and interpretability of the Euler-toured path, returning verbalized path as $P_{v}$. We return two different verbalized paths $P_{v}(\!q,\!P)\!=\!\textit{Verbalizer}(\!q,\!P)$ and $P_{\!v\!}(P)\!=\!\textit{Verbalizer}(\!P)$ where $q$ is the query node whose definition is used by the verbalizer, which is concatenated with path $P$ via the ([SEP]) token. The pre-trained encoder returns the following vectors,
@@ 4.4 Parent Verifier Module
In order to verify if the retrieved parent from LLM is the most appropriate one, existing works have employed discriminative verifiers Zeng et al. (2024a). However, these verifiers are often limited by their representation power. Moreover, if we use path-based verifier similar to TEMPORA, the verifier is again biased towards the ranked candidates, creating a bottleneck in verification. To overcome this, we instead utilize LLMs to reason over candidate paths within a batch to identify the most plausible hierarchical structure. However, direct reasoning over paths introduces instability, particularly when using smaller instruction-tuned LLMs. These models frequently generate inconsistent or malformed outputs, omitting full paths or introducing noise due to complex reasoning steps. To mitigate this, we compute the average token log-probability for each candidate path in the prompt and use these scores to rank them. This approach offers a more stable and interpretable signal for selecting the best-reasoned path.
The LLM is prompted with the hierarchical path of retrieved candidate from parent retriever module along with the path of remaining candidates within the batch to select the most appropriate path (see Appendix D for the prompt). Then, average token log probabilities are computed for all the paths in the batch to rank them to eliminate the possibility of the verifier selecting longer paths. The highest-scored candidate path is computed as,
@@ 5.1 Experimental Setup
We experiment on four open-source benchmarks: three from SemEval-2016 Task 13 Taxonomy Extraction Evaluation Bordea et al. (2016b), Science (SemEval-Sci or Sci), Environment (SemEval-Env or Env) and Food (SemEval-Food) while the other is a collection of 114 WordNet sub-taxonomies (with 10 to 50 nodes each) Bansal et al. (2014) (refer to Table 8 for statistics on benchmarks and Appendix G for train-test split and procedure to resolve definitions).
@@
Figure 3: Performance comparison of hierarchical and semantic reasoning on parent retrieval across all benchmarks. Here W/O means without.
Method Env Sci WordNet
Acc Wu&P Acc Wu&P Acc Wu&P
TEMPORA 67.31 82.99 64.71 87.42 49.78 84.58
Random Shuffle 5.76 47.64 2.35 48.95 1.51 38.95
@@ 5.2 Performance Comparison
Table 2 shows the performance of LORex on four benchmarks. As discussed in Appendices C and D, our framework uses prompts in an instruction-tuned paradigm. Therefore, we use three instruction-tuned LLMs , LLaMA-3.1-8B-Instruct (LORex8B-3.1I), LLaMA-3-8B-Instruct (LORex8B-3I), and LLaMA-2-7B-Chat (LORex7B-2C) to compare against state-of-the-art methods. As shown, LORex surpasses all first and second-generation discriminative and third-generation prompting methods, with LORex8B-3.1I achieving the best performance. Specifically, LORex achieves 12% improvement in accuracy and 5% improvement in Wu&P over the best extensively trained generative baseline, FLAME.
@@ 5.3 Ablation Studies
We conduct ablations to analyze every component of LORex. Specifically, we analyze (i) the impact of hierarchical and semantic reasoning on parent retrieval, (ii) effect of different types of reasoning verifiers, (iii) the influence of discriminative ranker, (iv) varying chunk sizes on the overall performance of LORex, and (v) impact of semantic filter module on the performance and latency of LORex.
Figure 4: Performance comparison of different types of verifiers on SemEval-2016 benchmark. "WoV" = without verifier, "CBV" denotes candidate-based verifier, while "PBV" denotes path-based verifier.
@@
We analyze the impact of hierarchical and semantic reasoning by utilizing paths and definitions, on parent retrieval in LORex, as shown in Fig. 3. Using both paths and definitions gives the highest performance while removing both components leads to the worst performance, dropping accuracy by 28.3% and Wu&P by 5.3%. Removing definitions reduces accuracy by 17.50% and Wu&P by 2.78%. Further, removing hierarchical paths exacerbates the performance degradation, leading to a 31.99% decrease in accuracy and a 9.78% reduction in Wu&P similarity scores.
@@
Fig. 4 compares LORex’s performance under different verification strategies. The path-based verifier (PBV) achieves the highest accuracy, outperforming the candidate-based verifier (CBV)  [^2] CBV refers to a verifier setup where the LLM is directly prompted to rank candidate terms without access to their hierarchical paths. The ranking is based solely on average log token probabilities over candidate names, without structured reasoning over taxonomy paths. by 47.04% in accuracy and 10.84% in Wu&P scores. Even without a verifier (WoV), retrieval remains 21.31% more accurate and 7.76% higher in Wu&P than CBV. The lowest performance is observed with CBV, where path information is absent, and verification relies solely on candidate evaluation, which introduces noise during verification and rejects relevant candidates. These results emphasize the importance of hierarchical reasoning during verification.
Impact of Discriminative Ranker LORex’s Performance.
Table 3 compares LORex’s performance with and without the use of a discriminative ranker. The discriminative ranker improves accuracy by 57.33% and Wu&P scores by 39.8%, highlighting its role in filtering relevant candidates and enhancing retrieval. In contrast, random shuffling of candidates disrupts the semantic ordering, often placing irrelevant or overly broad terms ahead of suitable hypernyms. This disordering hinders the retriever and verifier from progressing to later chunks where correct candidates may reside, thereby degrading overall system performance.
Setting Avg. Runs per Query Queries Resolved in $\leq$3 Runs for "Env"
Semantic Filter 1.7 41 / 52
W/O Semantic Filter 3.1 18 / 52
Table 4: Impact of the semantic filter module on reasoning efficiency. W/O denotes without.
Dataset Batches with True Parent Skipped Total Batches
Env 3 42
Sci 8 69
Food 13 238
WordNet 21 370
Table 5: Frequency of instances where the semantic filter skipped batches containing the true parent term.
Query Parent Retriever (Ground Truth) Path-Based Verifier Candidate-Based Verifier
@@
exhaustion of resources
a resource is a source or supply from which a benefit is produced and that has some utility. resource exhaustion refers to the depletion of these resources due to overuse or unsustainable practices.
natural resources (natural resources)
The most appropriate direct hypernym of ’exhaustion of resources’ is ’natural resources.’
exhaustion of resources > natural resources > resource > possession
The given path is ’exhaustion of resources -> natural resources -> resource -> possession’, which appropriately contextualizes the query.
resource
The query node ’exhaustion of resources’ is a type of resource.
@@
We analyze the impact of varying chunk sizes on LORex’s retrieval performance across SemEval-2016 benchmarks, as shown in Fig. 5. The results indicate that smaller chunk sizes (2–3) lead to performance degradation due to limited contextual information, while excessively large chunk sizes ($\geq$20) reduce effectiveness as the retriever encounters an increased number of irrelevant candidates, introducing noise in the chunk. The best performance is achieved with chunk sizes between 4 and 10, indicating a balance between relevant context depth and noise.
Impact of Semantic Filter Module on Performance and Latency of LORex.
As shown in Table 4, incorporating semantic filtering reduces the average number of retrieval-verification iterations per query from 3.1 to 1.7, and significantly increases the number of queries resolved within three runs. This demonstrates its utility in pruning irrelevant candidate batches early in the pipeline, reducing computational overhead and improving the latency of the framework. Furthermore, Table 5 shows that the semantic filter rarely eliminates batches containing the true parent node, only 3 out of 42 batches (for batch size=5) in the Environment dataset and similarly low rates across other datasets, indicating that the filter maintains high recall while enhancing precision. Failure cases of semantic filtering are discussed in Appendix K.